OpenStudy (anonymous):

What would $3,000 at 8.2% annual interest be if compounded monthly? How much would the investment be after 15 years? A. 9557.74 B. 9593.62 C. 9629.45 D. 9665.98

OpenStudy (anonymous):

Thanks for helping out Hugh ;)

OpenStudy (anonymous):

just trying to remember the formula.. 1 sec...

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

\[A = P(1 + (r/n)) ^(nt)\] something like that.. so P = 3000 r= 8.2 n = the number of compoundings per term, in this case 12 t = the term in this case years

OpenStudy (anonymous):

thats ^ nt

OpenStudy (anonymous):

couldnt get the formula to behave A = P(1 + (r/n))^(nt)

OpenStudy (anonymous):

A = 3000(1+(.082/12))^(12*15)

OpenStudy (anonymous):

Ithnink that's it.. run that calculation and you're good

OpenStudy (anonymous):

afk .. hgave to go